The position of a squirrel running in a park is given by π‘Ÿ = [(310.280π‘š/𝑠)𝑑 + (0.0360π‘š/𝑠 2 )𝑑 2 ]𝑖̂+ (0.0190π‘š/𝑠 3 )𝑑 3 𝑗̂. (a) What are 𝑣π‘₯(𝑑) and 𝑣𝑦(𝑑), the π‘₯- and 𝑦-components of the velocity of the squirrel, as functions of time? (b) At 𝑑 = 5.00 s, how far is the squirrel from its initial position? (c) At 𝑑 = 5.00 s, what are the magnitude and direction of the squirrel’s velocity?


1
Expert's answer
2021-09-20T10:01:21-0400

(a)

"v_x(t)=\dfrac{dr_x}{dt} = (0.28\ \dfrac{m}{s})+(0.072\ \dfrac{m}{s^2})t,""v_y(t)=\dfrac{dr_y}{dt} =(0.057\ \dfrac{m}{s^3})t^2."

(b) Let's first find "x"- and "y"-components of the final position of the squirrel at "t=5.0\ s":


"r_x=(0.28\ \dfrac{m}{s})\cdot5.0\ s+0.036\ \dfrac{m}{s^2}\cdot(5.0\ s)^2=2.3\ m,""r_y=(0.019\ \dfrac{m}{s^3})\cdot(5.0\ s)^3=2.4\ m."

Finally, we can find the final position of the squirrel at "t=5.0\ s" from the Pythagorean theorem:


"r=\sqrt{r_x^2+r_y^2}=\sqrt{(2.3\ m)^2+(2.4\ m)^2}=3.3\ m."

(c) Let's first find "x"- and "y"-components of the squirrel's velocity at "t=5.0\ s":


"v_x(t=5.0\ s)= (0.28\ \dfrac{m}{s})+(0.072\ \dfrac{m}{s^2})\cdot5.0\ s=0.64\ \dfrac{m}{s},""v_y(t=5.0\ s)=(0.057\ \dfrac{m}{s^3})\cdot(5.0\ s)^2=1.43\ \dfrac{m}{s}."

We can find the magnitude of the squirrel's velocity from the Pythagorean theorem:


"v=\sqrt{v_x^2+v_y^2}=\sqrt{(0.64\ \dfrac{m}{s})^2+(1.43\ \dfrac{m}{s})^2}=1.6\ \dfrac{m}{s}."

We can find the direction of the squirrel's velocity from the geometry:


"\theta=tan^{-1}(\dfrac{v_y}{v_x}),""\theta=tan^{-1}(\dfrac{1.43\ \dfrac{m}{s}}{0.64\ \dfrac{m}{s}})=66^{\circ}."